# Daco: Towards Application-Driven and Comprehensive Data Analysis via Code Generation

### v1

@@ 4 Experiments
The goal of our experiments is to verify that (1) augmenting language model with code generation can benefit data analysis, and (2) Daco-RL can further boost the answer helpfulness. To this purpose, we perform the following experiments:
Evaluated systems. We evaluate the code generation pipeline with ChatGPT and GPT-4. With the answer annotation generated by GPT-4, we further train a 6B CodeGeeX2-6B (Zheng et al., 2023a) model through both SFT and Daco-RL. For each of these models, we experiment with a baseline counterpart that does not include code generation and instead directly takes raw table content as input. We additionally experiment with two models specifically pre-trained on tabular data, TAPAS (Herzig et al., 2020) and TAPEX (Liu et al., 2022). TAPAS is a BERT-style model pre-trained to select relevant information from a table based on user query. For our dataset, we first use TAPAS to select relevant information and then use ChatGPT to interpret the selected information. TAPEX is a pre-trained encoder-to-decoder model. We fine-tune TAPEX with GPT-4-generated annotations.
Evaluation. The main metric we use is pairwise comparison of helpfulness as in Section 2. We use both ChatGPT and trained human annotators for the evaluation. We additionally report BLEU score, entailment score, and helpfulness evaluation for each individual bullet point. These metrics cannot holistically measure the analysis helpfulness, but can provide complementary insights for analyzing model performance. For entailment, we use an off-the-shelf NLI model to compute the probability that the model generation is entailed by the annotation. For point-wise evaluation, we ask the annotator to assign a score chosen from 0, 1 and 2 to each bullet point using the same standard as in human refinement of test set. Our human annotation achieves high agreement of 0.62 Cohen’s kappa for pairwise comparison of helpfulness, and 0.65 Cohen’s kappa for point-wise helpfulness evaluation.
@@ 4.1 Results
| TableQA Baselines | TAPAS | 337M | ✗ | 25.00 | 1.96 | 11.62 | 24.50 | 3.67 | 9.73 |
| TAPEX | 406M | ✗ | 14.79 | 3.34 | 14.60 | 6.00 | 3.50 | 13.81 | |
| Prompt- based LLM | ChatGPT | 20B | ✗ | 25.18 | 3.06 | 13.22 | 18.50 | 2.07 | 13.51 |
| GPT-4 | 175B | ✗ | 30.81 | 3.35 | 14.90 | 24.00 | 4.36 | 13.71 | |
| ChatGPT | 20B | ✓ | 35.74 | 2.74 | 14.22 | 27.27 | 2.59 | 14.51 | |
| GPT-4 | 175B | ✓ | 52.00 | 4.59 | 17.77 | 41.88 | 3.26 | 17.54 | |
| Finetuned LLM | SFT | 6B | ✗ | 21.51 | 2.30 | 14.47 | 9.50 | 2.65 | 13.63 |
| SFT | 6B | ✓ | 20.95 | 2.15 | 14.88 | 11.54 | 4.47 | 14.60 | |
| Daco-RL | 6B | ✓ | 28.54 | 3.65 | 13.13 | 21.05 | 5.98 | 11.80 | |
| Answer | Code | |||||||
| Help. | Help.-SFT | Entail. | BLEU | Info.-SFT | # lines | # API | Error | |
| SFT | 11.54 | 50.00 | 4.47 | 14.60 | 50.00 | 41.17 | 134 | 3.08 |
| Daco-RL | 21.05 | 58.49 | 5.98 | 11.80 | 57.86 | 39.66 | 145 | 2.75 |
| w/o | 18.75 | 51.92 | 3.21 | 13.17 | 52.56 | 32.54 | 120 | 3.84 |
| w/o | 8.79 | 40.00 | 3.13 | 11.46 | 44.57 | 14.30 | 108 | 3.91 |
@@
The main results are in Table 2. We have the following observations:
Code generation significantly helps data analysis, especially for zero-shot LLMs. ChatGPT and GPT-4 both enjoy a significant gain in most metrics, especially helpfulness, when equipped with code generation. As in Table 5, human evaluation further shows that GPT-4 with code generation has a significant 66.41 win rate over GPT-4 w/o code generation. After SFT, code generation brings less significant improvements, because SFT w/o code generation can simulate the behavior in GPT annotations and achieve competitive helpfulness. However, its mathematical and logical reasoning are not supported by code generation, so it produces more hallucination as reflected by the low entailment score on Test.
Our SFT model learns reasonable data analysis capabilities. By simulating GPT-4 behaviors, SFT with code generation achieves a reasonable helpfulness score and outperforms the TAPEX baseline, but still falls short compared to ChatGPT. We also evaluate the error rate (%) of generated code per step. Our SFT model has an error rate of 3.08%, which is reasonable low, but still much higher than ChatGPT (0.495%) and GPT-4 (0.491%).
@@
Daco-RL significantly improves over SFT. As shown in Table 2 and 3, Daco-RL significantly boosts the performance. Despite the difference in model size, DACO-RL outperforms ChatGPT w/o code generation on helpfulness and entailment metrics. When matched in size, DACO-RL significantly outperforms SFT by 7 points on helpfulness, further demonstrating its benefits. Human evaluation demonstrates a 57.72 win rate of Daco-RL over SFT as in Table 5. Our qualitative analysis shows that Daco-RL better focuses on user query, while SFT tends to display generic statistics that are less relevant to user query. An example is shown in Figure 6.
We further perform ablation study and report the results in Table 3. To directly compare the ablation models against SFT, we report the win rate of each model over SFT model. To evaluate the quality of code generation, we use ChatGPT to compare the “informativeness” of generated code against SFT, where informativeness refers to producing informative and insightful code execution outputs while staying relevant to the user query. We report a few additional statistics including the number of code lines and the number of different API calls. We observe that without our proposed two reward models (Daco-RL w/o ), using only answer RM significantly hurts the model generation, leading to short and less diverse code generation and thus less helpful final answers. Contribution RM and regularization RM encourage more diverse code generation and more helpful final answer production.
@@
Contribution RM favors API calls that extracts important information from tabular data but is also vulnerable to reward hacking. We report the Pearson correlation between API occurrence and contribution RM scores in Table 5. The functions rewarded most are related to extracting significant features (nlargest, sort_values), aggregating results (mean), and displaying specific information (print). In contrast, the least rewarded functions involve displaying generic statistics (describe) and wrangling data (merge, to_datetime, is_null) since they cannot directly contribute to the user query. Examples of generated code and their contribution RM scores are shown in Figure 6(a) and 6(b). However, we notice the concerningly high correlation between print function and contribution RM scores, which indicates the RL policy may exploit the correlation to hack reward. Figure 6(c) shows a typical reward hacking case, where the model achieves a high contribution RM score by printing. Our regularization RM learns to discourage such behavior and helps fix the gap.
@@
Evaluation on external test sets. To further analyze the effectiveness of Daco-RL, we evaluate our SFT and Daco-RL models on two external test sets: (1) data analysis benchmark InfiAgent-DA (Hu et al., 2024) (a concurrent work of ours), and (2) free-form table question answering dataset FeTaQA (Nan et al., 2022). We find that Daco-RL improves the accuracy over SFT on InfiAgent-DA (14.61 v.s. 12.92), especially over questions about summary statistics (14.86 v.s. 10.80) and correlation analysis (21.57 v.s. 14.86), which aligns with our evaluation results on Daco-RL dataset. On FeTaQA, Daco-RL retains similar performance (6.35 Rouge-L, 80.74 BERTScore) compared to SFT (6.39 Rouge-L, 80.68 BERTScore) since Daco-RL is not specifically trained to enhance information lookup capabilities.
@@ 5 Related Work
Table Analysis. Early work in table question answering (table QA) targets simple questions that requires table lookup and cell aggregations (Pasupat & Liang, 2015; Zhong et al., 2017; Iyyer et al., 2017; Yu et al., 2018; Nan et al., 2022). Later benchmarks further require free-form answer generation (Nan et al., 2022), multi-hop reasoning (Chen et al., 2021a; 2020) and mathematical reasoning (Zhu et al., 2021; Chen et al., 2021b; Lu et al., 2023). Despite the similar formulation between our task and existing table QA work, their focus are different: most existing table QA datasets focus on obtaining specific information, our data analysis queries can be complex and requires query decomposition and reasoning. Some concurrent work further targets comprehensive table analysis such as correlation analysis and causal reasoning (Nan et al., 2023; Hu et al., 2024; Liu et al., 2024). The main difference between this work to the concurrent work is our focus on addressing application-driven user queries.
Code Generation. Code generation benchmarks have been proposed for general-purpose programming (Austin et al., 2021; Hendrycks et al., 2021), math problems (Austin et al., 2021), and data science scenario (Lai et al., 2022; Huang et al., 2022). Similar to our work, some recent work allows the language model to interact with a code execution environment and receive execution outputs as feedback (Yang et al., 2023; Wang et al., 2023). The most relevant work is Cheng et al. (2023) that also addresses data analysis via code generation. Given a data analysis query, they use GPT-4 to first generate code and then provide an interpretation of the execution results. While their analysis queries are still relatively simple, this is an early exploration aiming at automating data analysis.
RLHF. Reinforcement learning from human feedback (RLHF) aims to optimize a language model against human preference (Ouyang et al., 2022; Touvron et al., 2023; Bai et al., 2022a; b; Ziegler et al., 2019; Wu et al., 2023). While traditionally RLHF uses a holistic reward score for the entire generation (Ziegler et al., 2019; Ouyang et al., 2022), recent work shows that dense reward scores for the intermediate reasoning steps are better learning signals (Lightman et al., 2023; Wu et al., 2023). These work uses expensive human annotation to collect annotations for the dense reward data. Compared to human preference, heuristic rewards are more accessible but may not align well with true reward. This gap can lead to reward hacking (Skalse et al., 2022; Pan et al., 2022). A common remedy is to use manually designed heuristics to penalize behaviors that potentially harm the true reward (Ouyang et al., 2022; Laidlaw et al., 2023). In this work, we train the regularizatoin RM to discourage reward hacking.
@@ 6 Conclusion
In this work, we propose a novel and challenging data analysis task, which involves decomposing user query into multiple perspectives, grounding each perspective to the input data and performing logical and mathematical reasoning. To support this task, we build the Daco dataset containing large-scale annotations automatically generated by GPT-4 and a small but high-quality test set with human curated annotations. We employ LLM enhanced with code generation to this task and evaluate three models on our dataset: zero-shot ChatGPT, zero-shot GPT-4 and a 6B SFT model. While GPT-4 consistently performs the best, SFT achieves reasonably good helpfulness with much less computation. On top of the SFT model, we further proposed our Daco-RL algorithm that significantly boosts the human evaluated helpfulness.
Appendix
Appendix A Additional Analysis of Daco Dataset
We perform additional analysis to verify the quality of our Daco dataset. We assess data quality based on comprehensiveness and agreement among annotators, which are two of the most commonly considered factors (Ehrlinger & Wöß, 2022).
We first measure the overlap between input queries to verify the diversity of automatically generated input queries. We compute the overlap among multiple queries over the same database using cosine similarity of Sentence-BERT embeddings. We use the all-MiniLM-L6-v2 model. We find that 46% pairs of generated queries have large difference. A small portion (2%) are repetitive with high similarity; since the percentage is small, it should not seriously affect the dataset quality. Details and qualitative examples are shown in Table 6.
Example |
| Large difference (similarity0.5) | 46 |
As a weather forecaster, I want to study the correlation between weather conditions and bike rentals. v.s. |
As a tourist attraction planner, I want to evaluate the bike-sharing program’s impact on tourism and visitor satisfaction. |
Similarity = 0.42 |
| Medium difference (0.5similarity0.8) | 52 |
As a farmer, I want to determine the suitable fruit varieties to grow on my farm. v.s. |
As a fruit exporter, I want to identify the fruits that meet export standards and have a longer shelf life. |
Similarity = 0.69 |
| Repetitive (similarity0.8) | 2 |
As a consultant for honey market, I want to study the honey production trend to recommend business strategies for my clients. v.s. |
As a curious analyst, I want to study the production trend to understand the US honey industry. |
Similarity = 0.85 |
We further evaluate the comprehensiveness of input queries, i.e. how many data columns are covered by the analysis. We apply heuristic rules to measure . On average, each analysis covers 71% data columns in the corresponding database. Among all data columns, 90% are covered by at least one data point. This verifies that our dataset achieves good coverage of the database columns.
Regarding agreement among annotators, as mentioned in the main content, the machine-generated queries are filtered by human annotators with a 0.62 Cohen’s kappa score, and our manual refinement of the test set also achieves a substantial 0.67 Cohen’s kappa score. These also verify the quality of our Daco dataset.
Appendix B Implementation Details

### v2

@@ 5 Experiments
5.1 Experimental Settings
Table QA baselines. We experiment with two models designed for table QA tasks, TAPAS [13] and TAPEX [19]. TAPAS is a BERT-style model pre-trained to select relevant information from a table based on user query. For our dataset, we first use TAPAS to select relevant information and then use ChatGPT to interpret the selected information. TAPEX is a encoder-to-decoder model pre-trained on table and SQL data. We fine-tune TAPEX with GPT-4-generated annotations.
Prompt-based LLMs. We evaluate the performance of ChatGPT and GPT-4 when enhanced by code generation. For comparison, we also experiment with a baseline counterpart that does not include code generation and instead directly takes raw table content as input.
Fine-tuned LLMs. With the answer annotation generated by GPT-4, we train a 6B CodeGeeX2-6B [44] model through SFT, RLHF, and fine-grained RLHF (denoted as FG-RLHF) [38].
We begin by training the model with SFT, using either all annotations including code generation or purely the final answer annotations. The former SFT model is further refined through RLHF with an end goal of optimizing the final answer ’s helpfulness. We model the helpness of final answer with an answer RM . Concretely, is trained on pairwise comparison data of output bullet points annotated by ChatGPT. To encourage diversity, we additionally impose repetition penalty by subtracting a similarity score between different bullet points from the final reward.
However, we observe that RLHF struggles to provide useful supervision signal to intermediate code generation since its reward is only applied to the final answer. Recent work has shown that providing dense reward for the intermediate steps can alleviate this issue [18, 38]. Thus, we propose to use fine-grained RLHF [38], where we introduce two novel reward models to provide dense reward signals for code generation: contribution RM and regularization RM . encourages code generation that better contributes to the final answer at each step, while helps prevent reward hacking. Concretely, we compute the similarity between final answer and the -th step observation to measure the contribution of generated code action . Using , we can rank the contribution of different steps, and then use the comparison pairs to train . However, this heuristic training of may lead to the well-studied reward misspecification issue termed reward hacking [34]. To mitigate this issue, we propose to regularize such behavior with . Given the misspecified reward model , we first train an RL model until its generations start to collapse to certain patterns, then we train to detect such patterns and thus assign lower scores.
Evaluation. The main metric we use is pairwise comparison of helpfulness as in Section 2. We use three LLM evaluators, GPT-4o mini, Claude 3.5 Sonnet, and Llama 3 8B [2], as well as trained human annotators for the evaluation. We additionally report BLEU score, entailment score, and helpfulness evaluation for each individual bullet point. These metrics cannot holistically measure the analysis helpfulness, but can provide complementary insights for analyzing model performance. For entailment, we use an off-the-shelf NLI model to compute the probability that the model generation is entailed by the annotation. For point-wise evaluation, we ask the annotator to assign a score chosen from 0 (not helpful), 1 (borderline helpful) and 2 (very helpful) to each bullet point using the same standard as in human refinement of test set. Our human annotation achieves high agreement of 0.62 Cohen’s kappa for pairwise comparison of helpfulness, and 0.65 Cohen’s kappa for point-wise helpfulness evaluation.
@@ 5.2 Results
| TableQA Baselines | TAPAS | 337M | ✗ | 19.19 | 1.96 | 11.62 | 16.50 | 3.67 | 9.73 |
| TAPEX | 406M | ✗ | 15.08 | 3.34 | 14.60 | 9.00 | 3.50 | 13.81 | |
| Prompt- based LLMs | ChatGPT | 20B† | ✗ | 19.31 | 3.06 | 13.22 | 13.50 | 2.07 | 13.51 |
| GPT-4 | 175B† | ✗ | 30.43 | 3.35 | 14.90 | 20.50 | 4.36 | 13.71 | |
| ChatGPT | 20B† | ✓ | 26.51 | 2.74 | 14.22 | 21.38 | 2.59 | 14.51 | |
| GPT-4 | 175B† | ✓ | 50.79 | 4.59 | 17.77 | 43.92 | 3.26 | 17.54 | |
| Finetuned LLMs | SFT | 6B | ✗ | 18.96 | 2.30 | 14.47 | 11.33 | 2.65 | 13.63 |
| SFT | 6B | ✓ | 13.73 | 2.15 | 14.88 | 9.83 | 4.47 | 14.60 | |
| RLHF | 6B | ✓ | 10.64 | 3.18 | 12.66 | 7.51 | 3.13 | 11.46 | |
| FG-RLHF | 6B | ✓ | 19.42 | 3.65 | 13.13 | 12.50 | 5.98 | 11.80 | |
@@
The main results are in Table 3. The key conclusions are as follows:
Proprietary LLMs demonstrate strong data analysis capabilities but still lag behind human performance. Proprietary LLMs, especially GPT-4, perform the best and significantly outperform other models on almost all metrics, particularly when enhanced with code generation. However, pairwise comparisons between model generation and human-refined annotations (TestH) show that even the best model only wins 41.88% of the time, indicating a gap in generating helpful data analysis.
@@
The fine-tuned models show promising data analysis capabilities, demonstrating the usefulness of our weak supervision data. By training with automatically generated annotations, SFT with code generation achieves a reasonable helpfulness score and outperforms the TAPEX baseline. Though RLHF negatively affects performance due to sparsity of reward signals, FG-RLHF with our dense reward models significantly improves the performance, outperforming SFT by 7 points in helpfulness. Despite the difference in model size, FG-RLHF outperforms ChatGPT without code generation in helpfulness and entailment metrics. Human evaluation shows a 57.72% win rate of FG-RLHF over SFT, as seen in Table 5. Our qualitative analysis indicates that FG-RLHF better focuses on user queries, while SFT tends to display generic statistics less relevant to user queries.
The fine-tuned models show promising data analysis capabilities, demonstrating the usefulness of our weak supervision data. By training with automatically generated annotations, SFT with code generation achieves a reasonable helpfulness score and outperforms the TAPEX baseline. Though RLHF negatively affect the performance due to the sparsity of reward signal, we see that FG-RLHF with our designed dense reward models significantly boosts the generation helpfulness, and outperforms SFT by 7 points on helpfulness, thus better aligning with human preference. Despite the difference in model size, FG-RLHF outperforms ChatGPT w/o code generation on helpfulness and entailment metrics. Human evaluation demonstrates a 57.72 win rate of FG-RLHF over SFT as in Table 5. Our qualitative analysis shows that FG-RLHF better focuses on user query, while SFT tends to display generic statistics that are less relevant to user query.
@@
We analyze the behavior of two reward models to better understand their effects on FG-RLHF. The contribution RM favors API calls that extract important information from tabular data but is vulnerable to reward hacking. As in Table 5, the functions rewarded most are related to extracting significant features (nlargest, sort_values), aggregating results (mean), and displaying specific information (print). In contrast, the least rewarded functions involve displaying generic statistics (describe) and wrangling data (merge, to_datetime, is_null) since they cannot directly contribute to the user query. However, the concerningly high correlation between print function and contribution RM scores indicates the policy may exploit the correlation to hack reward, which can be mitigated by employing the regularization RM.
@@
Evaluation on external test sets. We further evaluate our fine-tuned models on two external test sets: (1) InfiAgent-DA benchmark that focuses on complex but not application-driven data analysis [14], and (2) free-form table question answering dataset FeTaQA [26]. We find that FG-RLHF improves the accuracy over SFT on InfiAgent-DA (14.61 v.s. 12.92), especially over questions about summary statistics (14.86 v.s. 10.80) and correlation analysis (21.57 v.s. 14.86), which aligns with our evaluation results on FG-RLHF dataset. On FeTaQA, FG-RLHF retains similar performance (6.35 Rouge-L, 80.74 BERTScore) compared to SFT (6.39 Rouge-L, 80.68 BERTScore) since FG-RLHF is not specifically trained to enhance information lookup capabilities.
@@ 6 Related Work
Table Analysis. Early work in table question answering (table QA) targets simple questions that requires table lookup and cell aggregations [31, 46, 16, 42, 26]. Later benchmarks further require free-form answer generation [26], multi-hop reasoning [7, 8] and mathematical reasoning [47, 9, 24]. Despite the similar formulation between our task and existing table QA work, their focus are different: most existing table QA datasets focus on obtaining specific information, our data analysis queries can be complex and requires query decomposition and reasoning. Some concurrent work further targets comprehensive table analysis such as correlation analysis and causal reasoning [27, 14, 20]. The main difference between this work to the concurrent work is our focus on addressing application-driven and complex user queries, which requires more reasoning skills.
Code Generation. Code generation benchmarks have been proposed for general-purpose programming [3, 12], math problems [3], and data science scenario [17, 15]. Similar to our work, some recent work allows the language model to interact with a code execution environment and receive execution outputs as feedback [39, 37]. The most relevant work is [10] that also addresses data analysis via code generation. Given a data analysis query, they use GPT-4 to first generate code and then provide an interpretation of the execution results. While their analysis queries are still relatively simple, this is an early exploration aiming at automating data analysis.
LLM Agent. The notable capability of LLMs in reasoning and planning inspires many work to use them as intelligent agents for complex tasks like Minecraft gaming [36], robotics control [1], and web browsing [41]. In this work, our code generation pipeline can be considered as an adaptation of ReAct [40], where the LLM iteratively generates code as actions and reads execution results as observations. Although we have not yet explored more advanced agent designs, such as tool-crafting agents [6, 43] or self-reflection agents [33, 25], these approaches could be adapted to our task by redefining the action space as code generation and the observation space as execution results. We leave this exploration for future work.
@@ 7 Conclusion
In this work, we propose a novel and challenging data analysis task, which involves decomposing user query into multiple perspectives, grounding each perspective to the input data and performing logical and mathematical reasoning. To support this task, we build the Daco dataset containing large-scale annotations automatically generated by GPT-4 and a small but high-quality test set with human curated annotations. We evaluate three types models on our dataset: table QA models, prompt-based proprietary LLMs, and open LLMs fine-tuned on our automatically collected annotations. While GPT-4 consistently performs the best, the fine-tuned models achieves reasonably good helpfulness with much less computation. On top of the SFT model, we further show that fine-grained RLHF can be employed to boost helpfulness perceived by humans.
Acknowledgement
This work was partially supported by ByteDance during Xueqing’s internship with the company. This work was also supported partially by an Amazon gift grant.
Checklist
1.
For all authors…
Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
Did you describe the limitations of your work? [Yes] Section C in the supplementary materials
Did you discuss any potential negative societal impacts of your work? [N/A]
Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
2.
If you are including theoretical results…
Did you state the full set of assumptions of all theoretical results? [N/A]
Did you include complete proofs of all theoretical results? [N/A]
3.
If you ran experiments (e.g. for benchmarks)…
Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Code and data are released at https://github.com/shirley-wu/daco
Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Section 5 in the main content, and Section D in the supplementary materials
Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] Did not perform experiments of multiple random seeds due to resource constraints
Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] Section D in the supplementary materials
4.
If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
If your work uses existing assets, did you cite the creators? [Yes] Section 3 in the main content
Did you mention the license of the assets? [Yes] Spider [42] releases their dataset using Apache-2.0 license
Did you include any new assets either in the supplemental material or as a URL? [Yes] Dataset is released at https://github.com/shirley-wu/daco
Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] The annotators are trained and we have obtained their consent
Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] The data is public data and does not contain personally identifiable information or offensive content
5.
If you used crowdsourcing or conducted research with human subjects…
Did you include the full text of instructions given to participants and screenshots, if applicable? [No] Annotations are done through internal annotators and instructions are potentially confidential information
Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [No] Annotations are done through internal annotators and costs are potentially confidential information
Appendix A Appendix
Project website is at https://shirley-wu.github.io/daco/index.html. Data and code are released at https://github.com/shirley-wu/daco. Croissant metadata record is at https://github.com/shirley-wu/daco/blob/main/data/croissant.json. We license our resources under Apache-2.0 license.
We thereby state that we bear all responsibility in case of violation of rights, etc., and confirmation of the data license.
Appendix B Dataset Documentation
Below are dataset documentation following the framework from datasheets for datasets:
Motivation:
For what purpose was the dataset created? - For the novel task of data analysis as explained in the main content.
Who created the dataset and on behalf of which entity? - This dataset is created during a collaboration of ByteDance AI Lab and University of California, Los Angeles.
Who funded the creation of the dataset? - ByteDance
Composition:
What do the instances that comprise the dataset represent? - Each instance contains a database of tabular data, a question, a reasoning process including code snippets, and a final answer. Everything is in text format, except the database is stored as pd.DataFrame
How many instances are there in total? - As detailed in Table 1 in the main content.
Does the dataset contain all possible instances or is it a sample of instances from a larger set? - The dataset is not a sample from a larger set.
What data does each instance consist of? - Raw data.
Is there a label or target associated with each instance? - Yes, as explained in Section 3 in the main content.
Is any information missing from individual instances? - No
Are relationships between individual instances made explicit? - N/A
Are there recommended data splits? - Yes. We split the dataset randomly, and encourage people to follow this split for reproductivity. We also curate human annotations only for the test set.
Are there any errors, sources of noise, or redundancies in the dataset? - Yes. The input questions and answer annotations are generated by ChatGPT, which will inevitably contain errors. We try to manage the affect by manually filtering the questions, and by curating a test set of human refined answer annotations.
Is the dataset self-contained, or does it link to or otherwise rely on external resources? - Self-contained.
Does the dataset contain data that might be considered confidential? - No.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? - No.
Does the dataset identify any subpopulations? - No.
Is it possible to identify individuals, either directly or indirectly from the dataset? - No.
Does the dataset contain data that might be considered sensitive in any way? - No.
Collection process: as described in Section 3 in the main content.
Preprocessing/cleaning/labeling: we release the raw text data and do not perform any preprocessing/cleaning/labeling of the texts.
Uses:
Has the dataset been used for any tasks already? - The data analysis task, as in the main content
Is there a repository that links to any or all papers or systems that use the dataset? - https://github.com/shirley-wu/daco
What (other) tasks could the dataset be used for? - As in the main content
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses? - N/A
Are there tasks for which the dataset should not be used? - N/A
Distribution:
Will the dataset be distributed to third parties outside of the entity on behalf of which the dataset was created? - Yes
How will the dataset will be distributed? - https://github.com/shirley-wu/daco
When will the dataset be distributed? - Already released
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? - Yes, Apache-2.0 license
Have any third parties imposed IP-based or other restrictions on the data associated with the instances? - No
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? - No
Maintanance:
Who will be supporting/hosting/maintaining the dataset? - The dataset is not planned to be a dynamic dataset, but the authors will keep maintaining the github repo
How can the owner/curator/manager of the dataset be contacted? - Github or email xueqing.wu@cs.ucla.edu
Is there an erratum? - No
Will the dataset be updated? - No unless to correct errors
Will older versions of the dataset continue to be supported/hosted/maintained? - N/A
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? - Yes, please feel free to do that as long as citing our work and following the license
Appendix C Limitations
While we put forth to construct the first of its kind dataset, Daco, for the comprehensive data analysis task, the challenging nature of the data analysis process (which often requires certain domain expertise) itself presents two major limitations of this work: (1) It is expensive to build expert-annotated dataset. In our work, the large-scale annotations are automatically generated by GPT-4, and their quality cannot always be well guaranteed. Although we curate a test set refined by humans, those answers are initially generated by GPT-4, which may introduce biases to human annotators during the refinement process for the final answer. Annotations curated by experts from scratch might have higher quality, but they are indeed quite costly and can create other sort of alignment problems. (2) It is nontrivial to evaluate model generations. Evaluating the quality of the analyses is by itself challenging and requires data science expertise. For automatic evaluation, we use ChatGPT to rank the helpfulness of model generations, which can partially but not perfectly align with human preference. Additionally, ChatGPT cannot always robustly evaluate the correctness of model generations. We use an off-the-shelf NLI model to evaluate the entailment probability between human-refined ground truths and the model generations, which can partially reflect the correctness of model generations. However, the entailment probability prediction can sometime propagate errors which lead to false positives or negatives. We make efforts to alleviate such an issue by additionally collecting human evaluations, which are supposed to better reflect the answer quality, despite that humans can occasionally exhibit subjective evaluation patterns. Notice that our annotators do not fully check the correctness of the generated answers, where we task them to focus more on the helpfulness metrics defined in this work.
Appendix D Implementation Details